OpenStudy (anonymous):

Can someone please help me! \[3c^2/(c^2-2c) \div c/(c^2-4) \]

OpenStudy (tkhunny):

If it were me, I would factor everything that factors and then divide. Do you remember how to divide these sorts of things? (Reciprocal and multiply?)

OpenStudy (anonymous):

I dont remember anything thats why im in sumemr school :(

OpenStudy (anonymous):

(-1/c)(c^3+c^2-4c+8)

OpenStudy (anonymous):

if u turn the botom fraction upside down you can multiply accross with a common denominator

OpenStudy (anonymous):

So what would the answer be?

OpenStudy (anonymous):

i just wrote it above :)

OpenStudy (tkhunny):

Well, then you will fail again unless you get your act together. You WILL have to do better than that. \(c^{2} - 2c = c(c-2)\) This is the Distributive Property. Please factor \(c^{2}-4\). @groges, please read the Code of Conduct. Giving answer is not allowed. Giving bad answers is also discouraged. For example, there is no concern for common denominators when multiplying.

OpenStudy (anonymous):

Oh okay, sorry :)

OpenStudy (anonymous):

groges i msg you

OpenStudy (anonymous):

none of what you said is in the multiple choice

OpenStudy (tkhunny):

@groges That answer wit c^3 in the denominator is also incorrect. @Jessicax452 Getting someone else to do your homework is not a successful strategy.

OpenStudy (anonymous):

What is it u are asked to do Jessica? Is it to simplify?

OpenStudy (anonymous):

yeah

OpenStudy (tkhunny):

I'll give you one for free: \(\dfrac{3c^{2}}{c^{2} - 2c} / \dfrac{c}{c^{2} - 4} = \dfrac{3c^{2}}{c^{2} - 2c} \cdot \dfrac{c^{2} - 4}{c}\) This is the Reciprocal and Multiply concept. Now factor what can be factored \(\dfrac{3c^{2}}{c(c-2)} \cdot \dfrac{(c+2)(c-2)}{c}\) The first one, demonstrated before is the Distributive Property The second one, you are expected to recognize the "Difference of Squares". The first fraction has common factor in numerator and denominator \(\dfrac{3c}{c-2} \cdot \dfrac{(c+2)(c-2)}{c}\) Now for some commutative properties \(\dfrac{3c}{c}\cdot\dfrac{c-2}{c-2}\cdot (c+2)\) Some inverse and identity properties now gives \(3(c+2)\) You are expected to know ALL of this. There are more properties, too. You MUST get good at such manipulations or you will not do well. Time to get up to speed!

OpenStudy (anonymous):

who are you my fuuckin father?